Write an equation for each line.

1.) through (-2,1) and parallel to y = -3x+1

2.) through (-3,-1) and parallel to y = -2/5x-4

3.) through (-2,1) and horizontal

4.) through (-2,1) and vertical
1
Expert's answer
2010-09-30T05:07:08-0400
in the form y= mx+b, the slopes of parallel lines are equal to each other
1. y=-3x+b, It passes through the point (-2,1): 1=-3*(-2)+ b; b=-5.
thus the equation would be:
y= -3x - 5;

2. y=-2/5x+b, through (-3,-1): -1 = -2/5*(-3)+b ; b=-11/5.

y=-2/5x-11/5